# Posterior Label Smoothing for Node Classification

### v2

@@ Abstract
Label smoothing is a widely studied regularization technique in machine learning. However, its potential for node classification in graph-structured data, spanning homophilic to heterophilic graphs, remains largely unexplored. We introduce posterior label smoothing, a novel method for transductive node classification that derives soft labels from a posterior distribution conditioned on neighborhood labels. The likelihood and prior distributions are estimated from the global statistics of the graph structure, allowing our approach to adapt naturally to various graph properties. We evaluate our method on 10 benchmark datasets using eight baseline models, demonstrating consistent improvements in classification accuracy. The following analysis demonstrates that soft labels mitigate overfitting during training, leading to better generalization performance, and that pseudo-labeling effectively refines the global label statistics of the graph. Our code is available at https://github.com/ml-postech/PosteL.
@@ 1 Introduction
Soft label, which contains class-wise probabilities, has demonstrated remarkable success in training neural networks across various domains, including computer vision and natural language processing (lukov2022teaching; muller2019does; Szegedy_2016_CVPR; vasudeva2024geols; lsnlp_NIPS2017_3f5ee243; lsvision_21). One of the popular approaches to obtain a soft label is label smoothing, which introduces uniform noise into the ground-truth labels. Despite its simplicity, this technique effectively regularizes the output distribution and enhances generalization (pereyra2017regularizing). Knowledge distillation (kd_hinton_15) is another effective option, which trains a teacher model with a given one-hot label and utilizes its output as a soft-label to train the student model.
One of the most convincing explanations of why knowledge distillation works is that soft label enables the learning of “dark knowledge” included in instances (allen-zhu2023towards; kd_hinton_15). Since additional information captured by the teacher model that one-hot labels cannot convey is encoded as a soft label, the student model learns richer features.
@@
Existing approaches that generate soft labels in the graph domain build up the method based on a more specified assumption that nodes tend to share the same label with their neighboring nodes. Based on this assumption, they construct soft labels by naively aggregating the labels of neighboring nodes (wang2021structure; zhou2023adaptive). This approach aligns well with homophilic graphs, where nodes of the same class are likely to be connected, leading to improved generalization. However, it conflicts with the nature of heterophilic graphs, where edges frequently connect nodes with different labels.
@@ 2.1 Node Classification
Various works leverage graph structures in different ways to perform node classification. Early approaches such as GCN (kipf2016semi), GraphSAGE (sage2017_5dd9db5e), and GAT (velivckovic2017graph) aggregate neighbor representations under the homophilic assumption. For tackling class imbalance on homophilic graphs, GraphSMOTE (zhao2021graphsmote), ImGAGN (qu2021imgagn), and GraphENS (park2022graphens) have been proposed. Meanwhile, H2GCN (beyond_58ae23d8) and U-GCN (jin2021universal) enhance performance on heterophilic graphs by aggregating representations from multi-hop neighbors. Other research focuses on adaptively learning the graph structure itself. For instance, GPR-GNN (chien2020adaptive) and CPGNN (zhu2021graph) determine which nodes to aggregate, while ChebNet (defferrard2016convolutional), APPNP (gasteiger2018predict), and BernNet (he2021bernnet) focus on learning appropriate filters from graph signals.
@@ 2.2 Classification with Soft Labels
kd_hinton_15 demonstrate that training a small student model with soft labels derived from a large teacher model’s predictions outperforms training with one-hot labels. This approach, known as knowledge distillation (KD), has proven effective for both model compression and performance improvement (jiao-etal-2020-tinybert; vision_kd_2019; recommendation_kd_18).
Alternatively, simpler methods for generating soft labels exist. Label smoothing (Szegedy_2016_CVPR) adds uniform noise to one-hot labels, and its benefits have been widely explored. For instance, muller2019does show that the label smoothing improves model calibration, while lukasik2020does connect the label smoothing to label-correction techniques and demonstrate its utility in addressing label noise. The label smoothing is popular in computer vision (lukov2022teaching; vasudeva2024geols; lsvision_21) and NLP (guo2021label; song2020learning; lsnlp_NIPS2017_3f5ee243), yet it remains relatively underexplored in the graph domain.
To our knowledge, only two studies specifically propose label smoothing techniques for node classification. SALS (wang2021structure) smooths a node’s label to match those of its neighbors, and ALS (zhou2023adaptive) aggregates neighborhood labels with adaptive refinements. However, neither work focuses on heterophilic graphs, where nodes often connect to dissimilar neighbors. Meanwhile, other studies have proposed smoothing the prediction output based on the graph structure (xie2023label; zhang2021node), but their motivations differ substantially from the label smoothing approach investigated in this paper (e.g., they adjust output logits rather than training labels).
@@ 4 Theoretical Analysis of PosteL
Lemma 2 indicates that in a heterophilic graph, having fewer neighbors of label $k$ than of label 1k1-k can make the posterior favor $k$. Detailed proofs are provided in Appendix B. These two lemmas highlight the key difference between PosteL and the neighborhood aggregation method in wang2021structure. In their method, naively aggregating neighborhood labels for smoothing on heterophilic graphs results in the soft label being dominated by the majority neighborhood label. This majority dominance contradicts the inherent heterophilic property, where nodes are more likely to connect to dissimilar labels. In contrast, Lemma 2 demonstrates that PosteL assigns a lower probability to the majority neighborhood label in heterophilic graphs, thereby mitigating majority dominance, while Lemma 1 shows that PosteL effectively retains similarity to the majority neighbor label in homophilic graphs. Moreover, nodes in heterophilic graphs tend to connect with nodes that have different labels, which implies that $|\mathcal{N}_{y_{i}}(i)|<|\mathcal{N}_{1-y_{i}}(i)|$. In this case, PosteL preserves the ground-truth label as the most probable label. A similar effect is also found in homophilic graphs.
Figure 2: Toy example illustrating the difference between PosteL and SALS (wang2021structure). The leftmost column shows three examples of a target node (represented as T) with different local neighborhood structures. The second and third columns show how SALS and PosteL create soft labels with homophilic and heterophilic graphs, respectively.
@@
Table 1: Classification accuracy on 10 node classification datasets. $\Delta$ represents the performance improvement achieved by PosteL compared to the backbone model trained with the ground-truth label. All results of the backbone model trained with the ground-truth label are sourced from he2021bernnet.
@@
In Figure 2, we illustrate the difference between SALS (wang2021structure) and PosteL with different levels of homophily. The first column shows three examples of a target node (represented as T) with different local neighborhood structures. The second and third columns show how SALS and PosteL make soft labels with homophilic and heterophilic graphs, respectively. We visualize Lemma 3 through the heterophilic parts of the first and second graphs (rows) and Lemma 4 through the first and third graphs (rows). As the results indicate, both methods perform similarly in homophilic graphs but not in heterophilic ones. We note that the behavior of ALS (zhou2023adaptive) is challenging to analyze due to the presence of the learnable component in their method. Except for the learnable part, the basic aggregation method of ALS is similar to that of SALS.
@@ 5.1 Node Classification
We assess the performance of our method across 10 node classification datasets. To examine the effect of our method on diverse types of graphs, we conduct experiments on both homophilic and heterophilic graphs. For the homophilic setting, we evaluate our method on five datasets: Cora, CiteSeer, and PubMed, which are citation networks where nodes represent documents and edges correspond to citation links (sen2008collective; yang2016revisiting), as well as the Amazon co-purchase graphs Computers and Photo (mcauley2015image), where nodes represent products and edges indicate frequent co-purchases. For the heterophilic setting, we use five datasets: Chameleon and Squirrel, which are Wikipedia networks where nodes represent pages and edges correspond to mutual links (rozemberczki2021multi); Actor, a co-occurrence network where nodes represent actors and edges indicate co-appearances on the same Wikipedia pages (tang2009social); and Texas and Cornell, which are webpage graphs where nodes represent web pages and edges denote hyperlinks (pei2020geom). Detailed statistics of each dataset are illustrated in Appendix C.
@@
We evaluate the performance of PosteL across various backbone models, including a multi-layer perceptron (MLP) without graph structure, and seven widely used graph neural networks: GCN (kipf2016semi), GAT (velivckovic2017graph), APPNP (gasteiger2018predict), ChebNet (defferrard2016convolutional), GPR-GNN (chien2020adaptive), BernNet (he2021bernnet), and OrderedGNN (song2023ordered).
We follow the experimental setup and backbone implementations of he2021bernnet. Specifically, we use fixed 10 sets of train, validation, and test splits with ratios of 60%/20%/20%, respectively, and measure the accuracy at the lowest validation loss. Each model is trained for 1,000 epochs, with early stopping applied if the validation loss does not improve over the last 200 epochs. Details of the experimental setup, including the hyperparameter search spaces and additional implementation specifics, are provided in Appendix D.
We compare our method with two domain-agnostic soft labeling methods, including label smoothing (LS) (Szegedy_2016_CVPR) and knowledge distillation (KD) (kd_hinton_15), as well as two label smoothing methods designed for node classification: SALS (wang2021structure) and ALS (zhou2023adaptive).
@@ 5.2 Empirical Analysis
We hypothesize that correctly predicting PosteL labels, which encode local neighborhood information, enhances the model’s understanding of the graph structure and thereby improves generalization. Similar context-prediction strategies have been used as pretraining methods in previous studies (hu2019strategies; rong2020self). Loss curves for homophilic datasets are provided in Figure 11 and heterophilic in Figure 13 in Appendix G, showing consistent patterns across datasets.
@@ Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2024-00337955; RS-2023-00217286) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2024-00457882, National AI Research Lab Project; RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).
@@ D.1 Learning Hyperparameters
We largely follow the search space outlined in he2021bernnet:
@@ Appendix G Additional Experiments
We evaluated the runtime of PosteL on the ogbn-products dataset (hu2020open), which contains 2,449,029 nodes and 61,859,140 edges, to assess its computational efficiency on a large-scale graph. We measured the time excluding the training time for iterative pseudo-labeling. Generating soft labels with PosteL takes 5.65 seconds, whereas a single training epoch requires 19.11 seconds, indicating that PosteL can generate soft labels efficiently even on large-scale graphs.
@@
Figure 9 presents the t-SNE (van2008visualizing) plots of node embeddings from the GCN with the Chameleon and Squirrel datasets. The node color represents the label. For each dataset, the left plot visualizes the embeddings with the ground-truth labels, while the right plot visualizes the embeddings with PosteL labels. The visualization shows that the embeddings from the soft labels form tighter clusters compared to those trained with the ground-truth labels. This visualization results coincide with the t-SNE visualization of the previous work of muller2019does.
@@
We extend the experimental results presented in the main text to additional GNN backbones and datasets. Table 9 reports the classification accuracy for eight additional backbones, including APPNP (gasteiger2018predict), ChebNet (defferrard2016convolutional), MLP, GPR-GNN (chien2020adaptive), and OrderedGNN (song2023ordered). Figures 11 and 13 show training/validation/test loss curves comparing ground-truth labels, SALS (wang2021structure) labels, and PosteL labels on all 10 datasets used in our experiments. Figure 14 shows the estimated conditional distributions based on (i) training labels only, (ii) training labels combined with pseudo-labels, and (iii) all ground-truth labels.
@@
Table 9: Classification accuracy on 10 node classification datasets. $\Delta$ represents the performance improvement achieved by PosteL compared to the backbone model trained with the ground-truth label. All results of the backbone model trained with the ground-truth label are sourced from he2021bernnet.

### v3

@@ Abstract
Label smoothing is a widely studied regularization technique in machine learning. However, its potential for node classification in graph-structured data, spanning homophilic to heterophilic graphs, remains largely unexplored. We introduce posterior label smoothing, a novel method for transductive node classification that derives soft labels from a posterior distribution conditioned on neighborhood labels. The likelihood and prior distributions are estimated from the global statistics of the graph structure, allowing our approach to adapt naturally to various graph properties. We evaluate our method on 10 benchmark datasets using eight baseline models, demonstrating consistent improvements in classification accuracy. The following analysis demonstrates that soft labels mitigate overfitting during training, leading to better generalization performance, and that pseudo-labeling effectively refines the global label statistics of the graph.
Codehttps://github.com/ml-postech/PosteL
@@ 1 Introduction
Soft label, which contains class-wise probabilities, has demonstrated remarkable success in training neural networks across various domains, including computer vision and natural language processing (Lukov et al. 2022; Müller et al. 2019; Szegedy et al. 2016; Vasudeva et al. 2024; Vaswani et al. 2017; Zhang et al. 2021a). One of the popular approaches to obtain a soft label is label smoothing, which introduces uniform noise into the ground-truth labels. Despite its simplicity, this technique effectively regularizes the output distribution and enhances generalization (Pereyra et al. 2017). Knowledge distillation (Hinton et al. 2015) is another effective option, which trains a teacher model with a given one-hot label and utilizes its output as a soft-label to train the student model.
One of the most convincing explanations of why knowledge distillation works is that soft label enables the learning of “dark knowledge” included in instances (Allen-Zhu and Li 2023; Hinton et al. 2015). Since additional information captured by the teacher model that one-hot labels cannot convey is encoded as a soft label, the student model learns richer features.
@@
Existing approaches that generate soft labels in the graph domain build up the method based on a more specified assumption that nodes tend to share the same label with their neighboring nodes. Based on this assumption, they construct soft labels by naively aggregating the labels of neighboring nodes (Wang et al. 2021; Zhou et al. 2023). This approach aligns well with homophilic graphs, where nodes of the same class are likely to be connected, leading to improved generalization. However, it conflicts with the nature of heterophilic graphs, where edges frequently connect nodes with different labels.
@@ 2.1 Node Classification
Various works leverage graph structures in different ways to perform node classification. Early approaches such as GCN (Kipf and Welling 2016), GraphSAGE (Hamilton et al. 2017), and GAT (Veličković et al. 2017) aggregate neighbor representations under the homophilic assumption. For tackling class imbalance on homophilic graphs, GraphSMOTE (Zhao et al. 2021), ImGAGN (Qu et al. 2021), and GraphENS (Park et al. 2022) have been proposed. Meanwhile, H2GCN (Zhu et al. 2020) and U-GCN (Jin et al. 2021) enhance performance on heterophilic graphs by aggregating representations from multi-hop neighbors. Other research focuses on adaptively learning the graph structure itself. For instance, GPR-GNN (Chien et al. 2020) and CPGNN (Zhu et al. 2021) determine which nodes to aggregate, while ChebNet (Defferrard et al. 2016), APPNP (Gasteiger et al. 2018), and BernNet (He et al. 2021) focus on learning appropriate filters from graph signals.
@@ 2.2 Classification with Soft Labels
Hinton et al. (2015) demonstrate that training a small student model with soft labels derived from a large teacher model’s predictions outperforms training with one-hot labels. This approach, known as knowledge distillation (KD), has proven effective for both model compression and performance improvement (Jiao et al. 2020; Liu et al. 2019; Tang and Wang 2018).
Alternatively, simpler methods for generating soft labels exist. Label smoothing (Szegedy et al. 2016) adds uniform noise to one-hot labels, and its benefits have been widely explored. For instance, Müller et al. (2019) show that the label smoothing improves model calibration, while Lukasik et al. (2020) connect the label smoothing to label-correction techniques and demonstrate its utility in addressing label noise. The label smoothing is popular in computer vision (Lukov et al. 2022; Vasudeva et al. 2024; Zhang et al. 2021a) and NLP (Guo et al. 2021; Song et al. 2020; Vaswani et al. 2017), yet it remains relatively underexplored in the graph domain.
To our knowledge, only two studies specifically propose label smoothing techniques for node classification. SALS (Wang et al. 2021) smooths a node’s label to match those of its neighbors, and ALS (Zhou et al. 2023) aggregates neighborhood labels with adaptive refinements. However, neither work focuses on heterophilic graphs, where nodes often connect to dissimilar neighbors. Meanwhile, other studies have proposed smoothing the prediction output based on the graph structure (Xie et al. 2023; Zhang et al. 2021b), but their motivations differ substantially from the label smoothing approach investigated in this paper (e.g., they adjust output logits rather than training labels).
@@ 4 Theoretical Analysis of PosteL
Lemma 2 indicates that in a heterophilic graph, having fewer neighbors of label $k$ than of label 1k1-k can make the posterior favor $k$. Detailed proofs are provided in Appendix B. These two lemmas highlight the key difference between PosteL and the neighborhood aggregation method in Wang et al. (2021). In their method, naively aggregating neighborhood labels for smoothing on heterophilic graphs results in the soft label being dominated by the majority neighborhood label. This majority dominance contradicts the inherent heterophilic property, where nodes are more likely to connect to dissimilar labels. In contrast, Lemma 2 demonstrates that PosteL assigns a lower probability to the majority neighborhood label in heterophilic graphs, thereby mitigating majority dominance, while Lemma 1 shows that PosteL effectively retains similarity to the majority neighbor label in homophilic graphs. Moreover, nodes in heterophilic graphs tend to connect with nodes that have different labels, which implies that $|\mathcal{N}_{y_{i}}(i)|<|\mathcal{N}_{1-y_{i}}(i)|$. In this case, PosteL preserves the ground-truth label as the most probable label. A similar effect is also found in homophilic graphs.
Figure 2: Toy example illustrating the difference between PosteL and SALS (Wang et al. 2021). The leftmost column shows three examples of a target node (represented as T) with different local neighborhood structures. The second and third columns show how SALS and PosteL create soft labels with homophilic and heterophilic graphs, respectively.
@@
Table 1: Classification accuracy on 10 node classification datasets. $\Delta$ represents the performance improvement achieved by PosteL compared to the backbone model trained with the ground-truth label. All results of the backbone model trained with the ground-truth label are sourced from He et al. (2021).
@@
In Figure 2, we illustrate the difference between SALS (Wang et al. 2021) and PosteL with different levels of homophily. The first column shows three examples of a target node (represented as T) with different local neighborhood structures. The second and third columns show how SALS and PosteL make soft labels with homophilic and heterophilic graphs, respectively. We visualize Lemma 3 through the heterophilic parts of the first and second graphs (rows) and Lemma 4 through the first and third graphs (rows). As the results indicate, both methods perform similarly in homophilic graphs but not in heterophilic ones. We note that the behavior of ALS (Zhou et al. 2023) is challenging to analyze due to the presence of the learnable component in their method. Except for the learnable part, the basic aggregation method of ALS is similar to that of SALS.
@@ 5.1 Node Classification
We assess the performance of our method across 10 node classification datasets. To examine the effect of our method on diverse types of graphs, we conduct experiments on both homophilic and heterophilic graphs. For the homophilic setting, we evaluate our method on five datasets: Cora, CiteSeer, and PubMed, which are citation networks where nodes represent documents and edges correspond to citation links (Sen et al. 2008; Yang et al. 2016), as well as the Amazon co-purchase graphs Computers and Photo (McAuley et al. 2015), where nodes represent products and edges indicate frequent co-purchases. For the heterophilic setting, we use five datasets: Chameleon and Squirrel, which are Wikipedia networks where nodes represent pages and edges correspond to mutual links (Rozemberczki et al. 2021); Actor, a co-occurrence network where nodes represent actors and edges indicate co-appearances on the same Wikipedia pages (Tang et al. 2009); and Texas and Cornell, which are webpage graphs where nodes represent web pages and edges denote hyperlinks (Pei et al. 2020). Detailed statistics of each dataset are illustrated in Appendix C.
@@
We evaluate the performance of PosteL across various backbone models, including a multi-layer perceptron (MLP) without graph structure, and seven widely used graph neural networks: GCN (Kipf and Welling 2016), GAT (Veličković et al. 2017), APPNP (Gasteiger et al. 2018), ChebNet (Defferrard et al. 2016), GPR-GNN (Chien et al. 2020), BernNet (He et al. 2021), and OrderedGNN (Song et al. 2023).
We follow the experimental setup and backbone implementations of He et al. (2021). Specifically, we use fixed 10 sets of train, validation, and test splits with ratios of 60%/20%/20%, respectively, and measure the accuracy at the lowest validation loss. Each model is trained for 1,000 epochs, with early stopping applied if the validation loss does not improve over the last 200 epochs. Details of the experimental setup, including the hyperparameter search spaces and additional implementation specifics, are provided in Appendix D.
We compare our method with two domain-agnostic soft labeling methods, including label smoothing (LS) (Szegedy et al. 2016) and knowledge distillation (KD) (Hinton et al. 2015), as well as two label smoothing methods designed for node classification: SALS (Wang et al. 2021) and ALS (Zhou et al. 2023).
@@ 5.2 Empirical Analysis
We hypothesize that correctly predicting PosteL labels, which encode local neighborhood information, enhances the model’s understanding of the graph structure and thereby improves generalization. Similar context-prediction strategies have been used as pretraining methods in previous studies (Hu et al. 2019; Rong et al. 2020). Loss curves for homophilic datasets are provided in Figure 9 and heterophilic in Figure 10 in Appendix G, showing consistent patterns across datasets.
@@ Acknowledgements
We are grateful to Sangwoo Seo for providing insightful comments on this work. This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2024-00337955; RS-2023-00217286) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2024-00457882, National AI Research Lab Project; RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).
@@ D.1 Learning Hyperparameters
We largely follow the search space outlined in He et al. (2021):
@@ Appendix G Additional Experiments
We evaluated the runtime of PosteL on the ogbn-products dataset (Hu et al. 2020), which contains 2,449,029 nodes and 61,859,140 edges, to assess its computational efficiency on a large-scale graph. We measured the time excluding the training time for iterative pseudo-labeling. Generating soft labels with PosteL takes 5.65 seconds, whereas a single training epoch requires 19.11 seconds, indicating that PosteL can generate soft labels efficiently even on large-scale graphs.
@@
Figure 9 presents the t-SNE (Van der Maaten and Hinton 2008) plots of node embeddings from the GCN with the Chameleon and Squirrel datasets. The node color represents the label. For each dataset, the left plot visualizes the embeddings with the ground-truth labels, while the right plot visualizes the embeddings with PosteL labels. The visualization shows that the embeddings from the soft labels form tighter clusters compared to those trained with the ground-truth labels. This visualization results coincide with the t-SNE visualization of the previous work of Müller et al. (2019).
@@
We extend the experimental results presented in the main text to additional GNN backbones and datasets. Table 9 reports the classification accuracy for eight additional backbones, including APPNP (Gasteiger et al. 2018), ChebNet (Defferrard et al. 2016), MLP, GPR-GNN (Chien et al. 2020), and OrderedGNN (Song et al. 2023). Figures 11 and 13 show training/validation/test loss curves comparing ground-truth labels, SALS (Wang et al. 2021) labels, and PosteL labels on all 10 datasets used in our experiments. Figure 14 shows the estimated conditional distributions based on (i) training labels only, (ii) training labels combined with pseudo-labels, and (iii) all ground-truth labels.
@@
Table 9: Classification accuracy on 10 node classification datasets. $\Delta$ represents the performance improvement achieved by PosteL compared to the backbone model trained with the ground-truth label. All results of the backbone model trained with the ground-truth label are sourced from He et al. (2021).